which assembly of thermally equilibrated molecules has the more even distribution of molecules between the two lowest vibrational energy:a)O2 molecules at 300K or b)CO molecules at 1000K.
Assume that both molecules behave as harmonic oscillators, with the ground state vibrational frequencies for O2 and CO being 4.67x10^13 and 6.43x10^13 s^-1, respectively.
1
Expert's answer
2020-05-01T14:23:30-0400


For O2:

qvib = 2274

qvib / T = 2274 / 300 = 7.58 times = 8

For CO

qvib = 3087

qvib / T = 3087 / 1000 = 3.087 times = 3

8 > 3

Answer:

b)CO molecules at 1000K.